OpenStudy (anonymous):

helpp. Miles wants to purchase a home in six years. He will contribute $5000 each year to a savings account with 1.85% interest, compounded semiannually. What is the future value of this investment, when Maurice needs to make a down payment? $31,574.30 $31,866.37 $63,148.61 $63,732.73 @NaCl

OpenStudy (mathstudent55):

Are Maurice and Miles the same person?

OpenStudy (mathstudent55):

You need the compound interest formula.

OpenStudy (anonymous):

i have no idea lol i just noticed that.. can you help me

OpenStudy (mathstudent55):

Once again, you need the compound interest formula. Do you know it?

OpenStudy (anonymous):

no i dont.

OpenStudy (anonymous):

@NaCl

OpenStudy (mathstudent55):

\(F = A(1 + i) \left( \dfrac{(1 + i)^n - 1}{i} \right) \) where F = future value A = periodic payment i = interest rate per period written as decimal n = number of periods

OpenStudy (anonymous):

i cant solve it im not getting an answer choice

OpenStudy (mathstudent55):

\(F = A(1 + i) \left( \dfrac{(1 + i)^n - 1}{i} \right)\) \(F = $2500(1 + \dfrac{0.0185}{2} ) \left( \dfrac{(1 + \frac{0.0185}{2} )^{12} - 1}{\frac{0.0185}{2}} \right) \) \(F = $31,866.37 \)